# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ 3. Linear and bilinear Strichartz estimates
and this shows that we can essentially rely on (2), even if $b$ happens to be slighty smaller than $\frac{1}{2}$.
@@
just as in the proof of Theorem 1.1, it follows - taking into account the translation invariance of the Haar measure on $\mathbb{R}\times\mathbb{Z}$ that
@@
$\lVert I_{x}^{\frac{\alpha}{4}}P^{\alpha}(uv)\rVert_{L_{txy}^{2}}\lesssim\lVert J_{x}^{\frac{\alpha}{4}+\frac{1}{2}+}J_{y}^{\frac{1}{2}+}u\rVert_{X_{0,\frac{1}{4}+}}\lVert J_{x}^{\frac{\alpha}{4}+\frac{1}{2}+}v\rVert_{X_{0,\frac{1}{4}+}},$ (32)
@@
Let $T>0$, $0<\varepsilon\ll 1$ and $b>\frac{1}{2}$. Furthermore, if $u$, $v$, $w$ are time-localized [^3] Time localization is not actually required for the function $v$. functions with $J_{x}^{\frac{1}{5}}u\in X_{\varepsilon,b}$, $I_{y}^{\frac{1}{5}}v\in X_{\varepsilon,b}$, and $w\in X_{\frac{2}{15}+\varepsilon,b}$, then the following three estimates hold:
@@
$\lVert I_{x}^{\frac{1}{10}}v\rVert_{L_{txy}^{5}}\lesssim_{b}\lVert I_{y}^{\frac{1}{5}}v\rVert_{X_{\varepsilon,b}},$ (35)
@@
By interpolating the newly obtained $L^{4}$-estimate (2) with (12), we obtain (34), while (35) follows from interpolating (2) and (17). By decomposing on the Fourier side into the regions $A\coloneqq\{(\xi,q)\ |\ \lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}}\}$ and $A^{c}\coloneqq\{(\xi,q)\ |\ \lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}}\}$, estimate (36) follows from using estimate (34) on region $A$ and estimate (35) on region $A^{c}$. ∎
@@ 4.1. General LWP for gZK
Let $k\in\mathbb{N}^{\geq 2}$, and let $s_{0}(k)$ be as defined in Theorem 1.2. Then, for every $s(k)>s_{0}(k)$, there exists a sufficiently small number $\varepsilon=\varepsilon(s(k))>0$ such that the following estimate holds for all time-localized functions $u_{i}\in X_{s(k),\frac{1}{2}+\varepsilon}$:
$\lVert\partial_{x}\Bigg{(}\prod_{i=1}^{k+1}u_{i}\Bigg{)}\rVert_{X_{s(k),-\frac{1}{2}+2\varepsilon}}\lesssim\prod_{i=1}^{k+1}\lVert u_{i}\rVert_{X_{s(k),\frac{1}{2}+\varepsilon}}.$ (37)
@@
$I_{f}$ $\coloneqq\Bigg{\lvert}\int_{\mathbb{R}^{2}}\sum_{q\in\mathbb{Z}}\int_{\mathbb{R}^{2k}}\sum_{\begin{subarray}{c}q_{1},...,q_{k}\in\mathbb{Z}\\ \ast\end{subarray}}\xi\langle(\xi,q)\rangle^{s(k)}\Bigg{(}\prod_{i=1}^{k+1}\hat{u}_{i}(\tau_{i},\xi_{i},q_{i})\Bigg{)}\overline{\hat{f}}(\tau,\xi,q)\mathrm{d}\tilde{\tau}\mathrm{d}\tilde{\xi}\mathrm{d}(\tau,\xi)\Bigg{\rvert}$
$\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\prod_{i=1}^{k+1}\lVert u_{i}\rVert_{X_{s(k),\frac{1}{2}+\varepsilon}}$
@@
In this case, we have $\lvert(\xi_{i},q_{i})\rvert\lesssim 1$ and $\lvert(\xi,q)\rvert\lesssim 1$, for all $i\in\{1,2,3\}$, so that Sobolev’s embedding theorem applies without any derivative loss. Consequently, for any $0<\varepsilon\ll 1$ and $s\in\mathbb{R}$, we obtain
@@
This case can again be treated exactly as in (i.1) and (ii.1). For an arbitrary $s\in\mathbb{R}$ and $0<\varepsilon\ll 1$, we obtain
@@
In this final subcase, no further work is required, since we have $\lvert(\xi_{1},q_{1})\rvert\sim\lvert(\xi_{2},q_{2})\rvert\\ \sim\lvert(\xi_{3},q_{3})\rvert\sim\lvert(\xi_{4},q_{4})\rvert\sim\lvert(\xi_{i},q_{i})\rvert\sim\lvert(\xi_{5},q_{5})\rvert$, allowing us to argue as in case (iii.3.1). We apply the Airy $L^{6}$-estimate to the factor $u_{i}$, and use the optimized $L^{6}$-estimate for the factors $u_{1},...u_{4}$. For all remaining factors, we invoke the Sobolev embedding theorem, which altogether yields the desired estimate (37) for every $s>1-\frac{16}{9k}$. ∎
@@
The proof shows that local well-posedness for every $s>\frac{1}{2}$ in the case of the modified ZK equation can be achieved through the following simple alternative: for widely separated frequencies $\lvert(\xi_{i},q_{i})\rvert\gg\lvert(\xi_{j},q_{j})\rvert$, one applies the bilinear estimate (3), whereas for closely spaced frequencies $\lvert(\xi_{i},q_{i})\rvert\sim\lvert(\xi_{j},q_{j})\rvert$, one relies on the linear $L^{4}$-estimate (2). In the next section, we will incorporate the newly developed bilinear refinements (30) and the resonance function in order to further lower the well-posedness threshold to $s>\frac{3}{8}$.
@@ 4.2. Improved LWP for mZK
For every $s>\frac{3}{8}$, there exists a sufficiently small number $\varepsilon=\varepsilon(s)>0$ such that the following estimate holds for all time-localized functions $u_{i}\in X_{s,\frac{1}{2}+\varepsilon}$:
@@
for all $f\in X_{0,\frac{1}{2}-2\varepsilon}$ with $\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\leq 1$, and the convolution constraint being $\ast\coloneqq(\tau,\xi,q)=(\tau_{1}+\tau_{2}+\tau_{3},\xi_{1}+\xi_{2}+\xi_{3},q_{1}+q_{2}+q_{3})$. As in the proof of Proposition 4.1, we may assume without loss of generality that $\hat{u}_{i},\hat{f}\geq 0$ and $\lvert(\xi_{1},q_{1})\rvert\geq\lvert(\xi_{2},q_{2})\rvert\geq\lvert(\xi_{3},q_{3})\rvert$, where $(\xi_{i},q_{i})$ denotes the frequency variable associated with the factor $u_{i}$. Furthermore, we will arrange the $\xi_{i}$ for $i=1,...,3$ in ascending order according to the magnitude of their absolute values, and we will use the notation $\xi_{\text{min}}$, $\xi_{\text{med}}$, $\xi_{\text{max}}$ for this purpose. Lastly, we will also use $(\xi_{0},q_{0})$ as an alternative notation for the frequency variable $(\xi,q)$. With these preliminary considerations in place, we now turn to a detailed case-by-case analysis.
@@
(ii.2) $\lvert\xi_{0}\rvert\gg 1$, $\lvert\lvert(\xi_{i},q_{i})\rvert^{2}-\lvert(\xi_{j},q_{j})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{3}{2}}$, for some $i,j\in\{0,1,2,3\}$ with $i\neq j$: Given that $\lvert(\xi_{1},q_{1})\rvert\sim\lvert(\xi_{2},q_{2})\rvert\sim\lvert(\xi_{3},q_{3})\rvert$, we may, without loss of generality distinguish between two cases. If we have $\lvert\lvert(\xi_{1},q_{1})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{3}{2}}$, it follows that
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\lvert(\xi_{1},q_{1})\rvert^{2}-\lvert(\xi_{2},q_{2})\rvert^{2}\rvert^{\frac{1}{2}}\langle(\xi_{1},q_{1})\rangle^{s}\langle(\xi_{2},q_{2})\rangle^{-\frac{1}{8}+}$
$\ \ \ \cdot\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{3},q_{3})\rangle^{\frac{3}{8}+},$
@@
$I_{f}$ $\lesssim\lVert MP(J^{s}u_{1},J^{-\frac{1}{8}+}u_{2})\rVert_{L_{txy}^{2}}\lVert J^{0-}fJ^{\frac{3}{8}+}\tilde{u}_{3}\rVert_{L_{txy}^{2}}$
$\leq\lVert MP(J^{s}u_{1},J^{-\frac{1}{8}+}u_{2})\rVert_{L_{txy}^{2}}\lVert J^{0-}f\rVert_{L_{txy}^{4-}}\lVert J^{\frac{3}{8}+}\tilde{u}_{3}\rVert_{L_{Txy}^{4+}}.$
@@
$\lesssim\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}},$
and by choosing $\varepsilon>0$ small enough, we obtain the desired estimate (42) for every $s>\frac{3}{8}$. On the other hand, if $\lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{1},q_{1})\rvert^{2}\rvert\gtrsim\lvert\xi_{0}\rvert^{\frac{3}{2}}$, it holds that
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{1},q_{1})\rvert^{2}\rvert^{\frac{1}{2}}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{1},q_{1})\rangle^{-\frac{1}{8}+}$
$\ \ \ \cdot\langle(\xi_{2},q_{2})\rangle^{\frac{s}{2}+\frac{3}{16}+}\langle(\xi_{3},q_{3})\rangle^{\frac{s}{2}+\frac{3}{16}+},$
@@
$I_{f}\lesssim\lVert MP(J^{0-}f,J^{-\frac{1}{8}+}\tilde{u}_{1})\rVert_{L_{txy}^{2}}\lVert J^{\frac{s}{2}+\frac{3}{16}+}u_{2}\rVert_{L_{txy}^{4}}\lVert J^{\frac{s}{2}+\frac{3}{16}+}u_{3}\rVert_{L_{txy}^{4}}.$
@@
... $\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{1}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{s}{2}+\frac{3}{16}+,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{s}{2}+\frac{3}{16}+,\frac{1}{2}+\varepsilon}}$
$\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{1}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{2}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{s,\frac{1}{2}+\varepsilon}},$
@@
$\frac{s}{2}+\frac{3}{16}+\leq s\Leftrightarrow\frac{3}{8}+\leq s,$
@@
(ii.3) $\lvert\xi_{0}\rvert\gg 1$, $\lvert\lvert(\xi_{i},q_{i})\rvert^{2}-\lvert(\xi_{j},q_{j})\rvert^{2}\rvert\ll\lvert\xi_{0}\rvert^{\frac{3}{2}}$, for all $i,j\in\{0,1,2,3\}$ with $i\neq j$:
For the remainder of this proof, we may assume without loss of generality that
$(\xi_{1},\xi_{2},\xi_{3})=(\xi_{\text{min}},\xi_{\text{med}},\xi_{\text{max}})$.
@@
In this situation, we have $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{1}+\xi_{3}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{2}\rvert$, and $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{3}\rvert$, so that we obtain
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim\lvert\xi_{1}+\xi_{3}\rvert^{\frac{1}{4}}\langle(\xi_{1},q_{1})\rangle^{\frac{3}{8}+}\langle(\xi_{3},q_{3})\rangle^{\frac{3}{8}+}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{2},q_{2})\rangle^{s-}.$
If we now have $\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert\gtrsim\lvert\xi_{1}+\xi_{3}\rvert$, then an application of Parseval’s identity and Hölder’s inequality yields
$I_{f}$ $\lesssim\lVert I_{x}^{\frac{1}{4}}P^{1}(J^{\frac{3}{8}+}u_{1}J^{\frac{3}{8}+}u_{3})\rVert_{L_{txy}^{2}}\lVert J^{0-}fJ^{s-}\tilde{u}_{2}\rVert_{L_{txy}^{2}}$
$\leq\lVert I_{x}^{\frac{1}{4}}P^{1}(J^{\frac{3}{8}+}u_{1}J^{\frac{3}{8}+}u_{3})\rVert_{L_{txy}^{2}}\lVert J^{0-}f\rVert_{L_{txy}^{4-}}\lVert J^{s-}\tilde{u}_{2}\rVert_{L_{Txy}^{4+}},$
and the first factor can be treated using the bilinear refinement (30) of (2). For the remaining two factors, we proceed exactly as in the proof of case (ii.2), and this leads us to
$\lesssim\lVert u_{1}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{2}\rVert_{X_{s,\frac{1}{2}+\varepsilon}},$
which is the desired estimate, provided $\varepsilon>0$ is chosen small enough. On the other hand, in the case $\lvert 3(\xi_{0}+\xi_{2})^{2}-(q_{0}+q_{2})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{2}\rvert$, we have
$\lvert\xi_{0}\rvert\langle(\xi_{0},q_{0})\rangle^{s}\lesssim\lvert\xi_{0}+\xi_{2}\rvert^{\frac{1}{4}}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{2},q_{2})\rangle^{\frac{3}{8}+}\langle(\xi_{1},q_{1})\rangle^{s-}\langle(\xi_{3},q_{3})\rangle^{\frac{3}{8}+},$
and proceeding as before, this leads to
$I_{f}$ $\lesssim\lVert I_{x}^{\frac{1}{4}}P^{1}(J^{0-}fJ^{\frac{3}{8}+}\tilde{u}_{2})\rVert_{L_{txy}^{2}}\lVert J^{s-}u_{1}J^{\frac{3}{8}+}u_{3}\rVert_{L_{txy}^{2}}$
$\leq\lVert I_{x}^{\frac{1}{4}}P^{1}(J^{0-}fJ^{\frac{3}{8}+}\tilde{u}_{2})\rVert_{L_{txy}^{2}}\lVert J^{s-}u_{1}\rVert_{L_{txy}^{4}}\lVert J^{\frac{3}{8}+}u_{3}\rVert_{L_{txy}^{4}}.$
The first factor is now to be treated using the variant (33) of the bilinear refinement (30), while the remaining two factors are estimated as in case (ii.2). We obtain
$\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\lVert u_{2}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}\lVert u_{3}\rVert_{X_{\frac{3}{8}+,\frac{1}{2}+\varepsilon}},$
which is the desired estimate (42). We may therefore assume

### v3

@@ 3. Linear and bilinear Strichartz estimates
and this shows that we can essentially rely on (2), even if $b$ happens to be slightly smaller than $\frac{1}{2}$.
@@
just as in the proof of Theorem 1.1, it follows - taking into account the translation invariance of the Haar measure on $\mathbb{R}\times\mathbb{Z}$ - that
@@
$\lVert I_{x}^{\frac{\alpha}{4}}P^{\alpha}(uv)\rVert_{L_{txy}^{2}}\lesssim\lVert J_{x}^{\frac{\alpha}{4}}u\rVert_{X_{1+,\frac{1}{4}+}}\lVert J_{x}^{\frac{\alpha}{4}}v\rVert_{X_{0,\frac{1}{4}+}},$ (32)
@@
Let $T>0$, $0<\varepsilon\ll 1$ and $b>\frac{1}{2}$. Furthermore, if $u$, $v$, $w$ are time-localized [^3] Time localization is not actually required for the function $v$. functions with $J_{x}^{\frac{1}{5}}u\in X_{\varepsilon,b}$, $J_{y}^{\frac{1}{5}}v\in X_{\varepsilon,b}$, and $w\in X_{\frac{2}{15}+\varepsilon,b}$, then the following three estimates hold:
@@
$\lVert I_{x}^{\frac{1}{10}}v\rVert_{L_{txy}^{5}}\lesssim_{b}\lVert J_{y}^{\frac{1}{5}}v\rVert_{X_{\varepsilon,b}},$ (35)
@@
By interpolating the newly obtained $L^{4}$-estimate (2) with (12), we obtain (34), while (35) follows from interpolating (2) and (17). By decomposing on the Fourier side into the regions $A\coloneqq\{(\xi,q)\ |\ \lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}}\}$ and $A^{c}\coloneqq\{(\xi,q)\ |\ \lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}}\}$, estimate (36) follows from using estimate (34) on region AA and estimate (35) on region $A^{c}$. ∎
@@ 4.1. General LWP for gZK
Let $k\in\mathbb{N}^{\geq 2}$, and let $s_{0}(k)$ be as defined in Theorem 1.2. Then, for every $s>s_{0}(k)$, there exists a sufficiently small number $\varepsilon=\varepsilon(s)>0$ such that the following estimate holds for all time-localized functions $u_{i}\in X_{s,\frac{1}{2}+\varepsilon}$:
$\lVert\partial_{x}\Bigg(\prod_{i=1}^{k+1}u_{i}\Bigg)\rVert_{X_{s,-\frac{1}{2}+2\varepsilon}}\lesssim\prod_{i=1}^{k+1}\lVert u_{i}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}.$ (37)
@@
$I_{f}$ $\coloneqq\Bigg\lvert\int_{\mathbb{R}^{2}}\sum_{q\in\mathbb{Z}}\int_{\mathbb{R}^{2k}}\sum_{\begin{subarray}{c}q_{1},...,q_{k}\in\mathbb{Z}\\ \ast\end{subarray}}\xi\langle(\xi,q)\rangle^{s}\Bigg(\prod_{i=1}^{k+1}\hat{u}_{i}(\tau_{i},\xi_{i},q_{i})\Bigg)\overline{\hat{f}}(\tau,\xi,q)\mathrm{d}\tilde{\tau}\mathrm{d}\tilde{\xi}\mathrm{d}(\tau,\xi)\Bigg\rvert$
$\lesssim\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\prod_{i=1}^{k+1}\lVert u_{i}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}$
@@
In this case, we have $\lvert(\xi_{i},q_{i})\rvert\lesssim 1$ and $\lvert(\xi,q)\rvert\lesssim 1$, for all $i\in\{1,2,3\}$, so that Sobolev’s embedding theorem applies without any derivative loss. Consequently, for any $0<\varepsilon\ll 1$ and $s\geq 0$, we obtain
@@
This case can again be treated exactly as in (i.1) and (ii.1). For an arbitrary $s\geq 0$ and $0<\varepsilon\ll 1$, we obtain
@@
In this final subcase, no further work is required, since we have $\lvert(\xi_{1},q_{1})\rvert\sim\lvert(\xi_{2},q_{2})\rvert\\ \sim\lvert(\xi_{3},q_{3})\rvert\sim\lvert(\xi_{4},q_{4})\rvert\sim\lvert(\xi_{i},q_{i})\rvert\sim\lvert(\xi_{5},q_{5})\rvert$, allowing us to argue as in case (iii.3.1). We apply the Airy $L^{6}$-estimate to the factor $u_{i}$, and use the optimized $L^{6}$-estimate for the factors $u_{1},...,u_{4}$. For all remaining factors, we invoke the Sobolev embedding theorem, which altogether yields the desired estimate (37) for every $s>1-\frac{16}{9k}$. ∎
@@
The proof shows that local well-posedness for every $s>\frac{1}{2}$ in the case of the modified ZK equation can be achieved through the following simple alternative: for widely separated frequencies $\lvert(\xi_{i},q_{i})\rvert\gg\lvert(\xi_{j},q_{j})\rvert$, one applies the bilinear estimate (3), whereas for closely spaced frequencies $\lvert(\xi_{i},q_{i})\rvert\sim\lvert(\xi_{j},q_{j})\rvert$, one relies on the linear $L^{4}$-estimate (2). In the next section, we will incorporate the newly developed bilinear refinements (30) and the resonance function in order to further lower the well-posedness threshold to $s>\frac{11}{24}$.
@@ 4.2. Improved LWP for mZK
For every $s>\frac{11}{24}$, there exists a sufficiently small number $\varepsilon=\varepsilon(s)>0$ such that the following estimate holds for all time-localized functions $u_{i}\in X_{s,\frac{1}{2}+\varepsilon}$:
@@
for all $f\in X_{0,\frac{1}{2}-2\varepsilon}$ with $\lVert f\rVert_{X_{0,\frac{1}{2}-2\varepsilon}}\leq 1$, and the convolution constraint being $\ast\coloneqq(\tau,\xi,q)=(\tau_{1}+\tau_{2}+\tau_{3},\xi_{1}+\xi_{2}+\xi_{3},q_{1}+q_{2}+q_{3})$. As in the proof of Proposition 4.1, we may assume without loss of generality that $\hat{u}_{i},\hat{f}\geq 0$ and $\lvert(\xi_{1},q_{1})\rvert\geq\lvert(\xi_{2},q_{2})\rvert\geq\lvert(\xi_{3},q_{3})\rvert$, where $(\xi_{i},q_{i})$ denotes the frequency variable associated with the factor $u_{i}$. Furthermore, we will arrange the $\xi_{i}$ for i=1,,3i=1,...,3 in ascending order according to the magnitude of their absolute values, and we will use the notation $\xi_{\text{min}}$, $\xi_{\text{med}}$, $\xi_{\text{max}}$ for this purpose. Lastly, we will also use $(\xi_{0},q_{0})$ as an alternative notation for the frequency variable $(\xi,q)$. With these preliminary considerations in place, we now turn to a detailed case-by-case analysis.
@@
(ii.2) ${\lvert}\xi_{0}{\rvert}\gg 1\text{, }{\lvert}{\lvert}(\xi_{i},q_{i}){\rvert}^{2}-{\lvert}(\xi_{j},q_{j}){\rvert}^{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{7}{6}}\text{, for some }i,j\in{\{0,1,2,3\}}\text{ with }i\neq j$:
Given that ${\lvert}(\xi_{1},q_{1}){\rvert}\sim{\lvert}(\xi_{2},q_{2}){\rvert}\sim{\lvert}(\xi_{3},q_{3}){\rvert}$, we may, without loss of generality distinguish between two cases. If we have ${\lvert}{\lvert}(\xi_{1},q_{1}){\rvert}^{2}-{\lvert}(\xi_{2},q_{2}){\rvert}^{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{7}{6}}$, it follows that
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}{\lvert}(\xi_{1},q_{1}){\rvert}^{2}-{\lvert}(\xi_{2},q_{2}){\rvert}^{2}{\rvert}^{\frac{1}{2}}\langle(\xi_{1},q_{1})\rangle^{s}\langle(\xi_{2},q_{2})\rangle^{-\frac{1}{24}+}$
$\ \ \ \cdot\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{3},q_{3})\rangle^{\frac{11}{24}+},$
@@
$I_{f}$ $\lesssim{\lVert}MP(J^{s}u_{1},J^{-\frac{1}{24}+}u_{2}){\rVert}_{L_{txy}^{2}}{\lVert}J^{0-}fJ^{\frac{11}{24}+}\tilde{u}_{3}{\rVert}_{L_{txy}^{2}}$
$\leq{\lVert}MP(J^{s}u_{1},J^{-\frac{1}{24}+}u_{2}){\rVert}_{L_{txy}^{2}}{\lVert}J^{0-}f{\rVert}_{L_{txy}^{4-}}{\lVert}J^{\frac{11}{24}+}\tilde{u}_{3}{\rVert}_{L_{Txy}^{4+}}.$
@@
$\lesssim{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}}{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}},$
and by choosing $\varepsilon>0$ small enough, we obtain the desired estimate (42) for every $s>\frac{11}{24}$. On the other hand, if ${\lvert}{\lvert}(\xi_{0},q_{0}){\rvert}^{2}-{\lvert}(\xi_{1},q_{1}){\rvert}^{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{7}{6}}$, it holds that
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}{\lvert}(\xi_{0},q_{0}){\rvert}^{2}-{\lvert}(\xi_{1},q_{1}){\rvert}^{2}{\rvert}^{\frac{1}{2}}\langle(\xi_{0},q_{0})\rangle^{0-}\langle(\xi_{1},q_{1})\rangle^{-\frac{1}{24}+}$
$\ \ \ \cdot\langle(\xi_{2},q_{2})\rangle^{\frac{s}{2}+\frac{11}{48}+}\langle(\xi_{3},q_{3})\rangle^{\frac{s}{2}+\frac{11}{48}+},$
@@
$I_{f}\lesssim{\lVert}MP(J^{0-}f,J^{-\frac{1}{24}+}\tilde{u}_{1}){\rVert}_{L_{txy}^{2}}{\lVert}J^{\frac{s}{2}+\frac{11}{48}+}u_{2}{\rVert}_{L_{txy}^{4}}{\lVert}J^{\frac{s}{2}+\frac{11}{48}+}u_{3}{\rVert}_{L_{txy}^{4}}.$
@@
... $\lesssim{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{1}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{\frac{s}{2}+\frac{11}{48}+,\frac{1}{2}+\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{\frac{s}{2}+\frac{11}{48}+,\frac{1}{2}+\varepsilon}}$
$\lesssim{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{1}{\rVert}_{X_{\frac{11}{24}+,\frac{1}{2}+\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}},$
@@
$\frac{s}{2}+\frac{11}{48}+\leq s\Leftrightarrow\frac{11}{24}+\leq s,$
@@
(ii.3) ${\lvert}\xi_{0}{\rvert}\gg 1\text{, }{\lvert}{\lvert}(\xi_{i},q_{i}){\rvert}^{2}-{\lvert}(\xi_{j},q_{j}){\rvert}^{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{7}{6}}\text{, for all }i,j\in{\{0,1,2,3\}}\text{ with }i\neq j$:
For the remainder of this proof, we may assume without loss of generality that $(\xi_{1},\xi_{2},\xi_{3})=(\xi_{\text{min}},\xi_{\text{med}},\xi_{\text{max}})$.
@@
In this situation, we have ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}$ and ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{2}+\xi_{3}{\rvert}$, so that we obtain
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}^{\frac{1}{4}}\langle(\xi_{1},q_{1})\rangle^{s-}\langle(\xi_{3},q_{3})\rangle^{\frac{1}{4}+}{\lvert}\xi_{0}+(-\xi_{2}){\rvert}^{\frac{1}{4}}\langle(\xi_{0},q_{0})\rangle^{0-}$
$\ \ \ \cdot\langle(-\xi_{2},-q_{2})\rangle^{\frac{1}{4}+}$
and
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}$ $\lesssim{\lvert}\xi_{2}+\xi_{3}{\rvert}^{\frac{1}{4}}\langle(\xi_{2},q_{2})\rangle^{s-}\langle(\xi_{3},q_{3})\rangle^{\frac{1}{4}+}{\lvert}\xi_{0}+(-\xi_{1}){\rvert}^{\frac{1}{4}}\langle(\xi_{0},q_{0})\rangle^{0-}$
$\ \ \ \cdot\langle(-\xi_{1},-q_{1})\rangle^{\frac{1}{4}+},$
respectively. If we now have ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\gtrsim{\lvert}\xi_{1}+\xi_{3}{\rvert}$, then an application of Parseval’s identity and Hölder’s inequality yields
$I_{f}\lesssim{\lVert}I_{x}^{\frac{1}{4}}P^{1}(J^{s-}u_{1}J^{\frac{1}{4}+}u_{3}){\rVert}_{L_{txy}^{2}}{\lVert}I_{x}^{\frac{1}{4}}P^{1}(J^{0-}fJ^{\frac{1}{4}+}\tilde{u}_{2}){\rVert}_{L_{txy}^{2}},$
and the first factor can be treated using the bilinear refinement (30) of (2), while the second factor can be handled using the variant (33) of (30). This leads us to
... $\lesssim{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}{\lVert}u_{3}{\rVert}_{X_{\frac{1}{4}+,\frac{1}{2}+\varepsilon}}{\lVert}f{\rVert}_{X_{0,\frac{1}{2}-2\varepsilon}}{\lVert}u_{2}{\rVert}_{X_{\frac{1}{4}+,\frac{1}{2}+\varepsilon}}$
with the last step being valid for all $s>\frac{1}{4}$, provided $\varepsilon>0$ is chosen small enough. In the case ${\lvert}3(\xi_{2}+\xi_{3})^{2}-(q_{2}+q_{3})^{2}{\rvert}\gtrsim{\lvert}\xi_{2}+\xi_{3}{\rvert}$, one uses the second pointwise estimate given above and follows the same argument to again reach $s>\frac{1}{4}$. We may therefore assume